# Mode-sum prescription for renormalized expectation values for a charged quantum scalar field on a charged black hole

### v1

@@ I Introduction
The development, over the past ten years, of two new methods for computing renormalized expectation values has greatly facilitated these calculations and enabled the properties of a neutral quantum scalar field on a much wider range of black hole space-times to be explored. The “pragmatic mode-sum” method [13, 14] works on the original Lorentzian space-time and is notable for its applicability to rotating as well as static black holes [15]. In contrast, the “extended coordinates” method [16] involves the Euclidean space-time. The key feature of the latter approach is that the Hadamard parametrix to any order in coordinate separation can be expressed as a mode-sum, resulting in a renormalized Green function that converges to any order we desire; the rate of convergence being directly linked to the order of the parametrix we subtract. This method also has the advantage that it is approximately uniform, including close to the horizon.
@@ II Euclidean Green function
$G_{\textrm{\tiny{E}}}(x,x^{\prime})=G^{\textrm{\tiny{(S)}}}(x,x^{\prime})+G^{\textrm{\tiny{(A)}}}(x,x^{\prime})$ (14a)
@@ III Expansion of the Hadamard Parametrix
and $\Delta r=r-r^{\prime}$ is the separation of the points in the radial direction. The coefficients in the expansions (22) can then be found by substituting (22) into the relevant equations (18, 21) and solving order by order, treating $\varpi\sim s\sim\Delta r\sim\mathcal{O}(\epsilon)$. A key feature of this particular choice of “extended coordinates” is that in the limit $\Delta r\to 0$, we have $s^{2}=2\sigma+\mathcal{O}(\epsilon^{3})$. Note also that, although the Hadamard parametrix itself (16) is independent of the quantum state, this particular choice of expansion parameters (23) is adapted specifically to quantum fields in thermal states. This is easily seen in the fact that the extended time coordinate $\varpi$ (23a) is periodic with periodicity related to $\kappa$. This is an essential feature of the method, since the goal is to obtain a mode-sum representation of the parametrix (16) that has the same form as the Euclidean Green function (II) so that a mode-by-mode subtraction can be performed.
@@
Let us deal first with the direct part of the parametrix, $U(x,x^{\prime})/\sigma(x,x^{\prime})$. If we expand the direct part in our extended coordinates up to order $\mathcal{O}(\epsilon^{2m})$ and take the partial coincidence limit $\Delta r=0$, we obtain
@@
The coefficients are now labelled by a $\mathcal{T}$ to indicate these come from the tail part of the parametrix and the superscript $(\mathrm{L})$ indicates coefficients of terms that contain a logarithm. Again, the first few of these coefficients are listed in Appendix A.
@@ IV.1 Euclidean vs Lorentzian expressions
An immediate difficulty arises when one comes to computing renormalized expectation values using Euclidean methods: how do we translate the formal Lorentzian expressions for the renormalized expectation values [6] into their Euclidean counterpart? For a neutral scalar field on a static spacetime, this is a straightforward matter since one simply replaces the Feynman Green function with the corresponding Euclidean Green function via the mapping
@@
$-iG_{\textrm{\tiny{F}}}(x,x^{\prime})=\langle 0|\mathrm{T}\{\hat{\Phi}(x)\,\hat{\Phi}^{\dagger}(x^{\prime})\}|0\rangle$ (29)
@@ IV.3 Renormalized current
Now the only nonzero component of the current for a charged scalar field on a static spacetime with an electrostatic potential is the time component, which is
@@
$j_{nl}^{(m)}(r)=$ $\sum_{a=0}^{m-1}\sum_{b=0}^{a}\sum_{k=0}^{m-a}\binom{\tfrac{1}{2}}{0}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{DO}_{ab}^{(\textrm{\tiny{R}})}(r)\Big{[}(2a+2b+1)\Psi_{nl}(a+k,b|r)-(2b+2)f(r)\Psi_{nl}(a+k,b+1|r)\Big{]}$
$+\sum_{a=1}^{m-2}\sum_{b=0}^{a-1}\sum_{k=0}^{m-a-2}\binom{\tfrac{1}{2}}{0}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{TO}_{ab}^{(\textrm{\tiny{R}})}(r)\Big{[}(2a+2b+3)\Psi_{nl}(a+k+1,b|r)$
@@
$+\sum_{a=0}^{m-2}\sum_{b=0}^{a}\sum_{k=0}^{m-a-2}\binom{\tfrac{1}{2}}{0}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{TO}_{ab}^{(\textrm{\tiny{L}})}(r)\Big{[}(2b+1)\chi_{nl}(a,b+k|r)+(2a-2b)f(r)\chi_{nl}(a-1,b+k+1|r)\Big{]}$
$+\sum_{a=0}^{m-2}\sum_{k=0}^{m-a-2}\binom{\tfrac{1}{2}}{0}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{TO}_{aa}^{(\textrm{\tiny{L}})}(r)2f(r)\Psi_{nl}(a+k+1,0|r)\Big{]}.$ (60b)
@@ IV.4 Renormalized stress energy tensor
using Synge’s Rule and the fact that $w$ is independent of time. The last term on the right-hand side we can compute by simply taking a radial derivative of the interpolation function for $w$. The salient point is that now the covariant derivatives in the first term on the right-hand side are at different spacetime points and so act as partial derivatives. This can now be computed from taking time derivatives of the mode-sum representation of $W^{\textrm{\tiny{(S)}}}(x,x^{\prime})$ already derived (IV.2). Explicitly, we have
@@
$=$ $~{}\frac{1}{8\pi^{2}}\sum_{l=0}^{\infty}(2l+1)\sum_{n=-\infty}^{\infty}\frac{n^{2}\kappa^{2}}{f}\left\{g_{nl}(r)-k_{nl}^{(m)}(r)\right\}$
@@
$\widetilde{w}^{\phi}{}_{\phi}=\widetilde{w}^{\theta}{}_{\theta}=-\Big{[}\nabla^{\phi^{\prime}}\nabla_{\phi}W^{\textrm{\tiny{(S)}}}(x,x^{\prime})\Big{]}+\frac{f}{2r}w_{,r},$ (72)
@@ V.2 Numerical results
We now use the expressions of the previous section (IV.2, 54, 68) to calculate the renormalized scalar condensate, current and RSET. We consider a Reissner-Nordström black hole with charge $Q=M/2$ and use units in which $M=1$. In these units the event horizon is located at $r=r_{+}=1+{\sqrt{3}}/2\approx 1.86602540$ to 9 s.f. (77). We assume that the scalar field is minimally coupled to the scalar curvature, so that $\xi=0$, and set the scalar field mass to be $\mu M=1/10$, whilst the scalar field charge is $qM=1/4$. With these parameters we have $\mu>\frac{qQ}{r_{+}}$ and therefore no classical charge superradiance [31]. In this case the Euclidean Green function (II) is uniquely defined. The presence of superradiant modes complicates the definition of a Hartle-Hawking-like state for a charged scalar field on a Reissner-Nordström black hole [20] and therefore, the construction of the Euclidean Green function. For this reason, we do not consider the possibility of charge superradiance in this paper, although we expect the framework presented in this paper for implementing Hadamard renormalization will be unaffected by the presence of superradiant modes. Finally, we also set the arbitrary renormalization length scale $L=M=1$ (16).
@@
The next stage is to compute the mode sums appearing in (IV.2, 60a, IV.4, IV.4), performing the sum over $n$ first, before the sum over $l$. As in the neutral case [23, 24], the summand in the sum relevant for calculating the scalar condensate (IV.2), namely $\{\Re\left(g_{nl}(r)\right)-k^{(m)}_{nl}(r)\}$, converges as ${\mathcal{O}}(n^{-2m-3})$ for fixed $l$ and large $n$, as demonstrated in Fig. 1 for $m=0,1,2$.
@@
The number of modes required for final answers for the renormalized expectation values to the desired precision can be reduced by increasing the order $m$ in the expansion of the Hadamard parametrix (IV.2, 60), at the cost of increased computation time of the required regularization parameters (42, 46). Table 1 shows the effect of increasing the expansion order $m$ and $l_{\rm{max}}$ on the convergence of results for the RSET component $\langle\hat{T}^{t}_{~{}t}\rangle$, with $n_{\rm{max}}$ fixed to be $n_{\rm{max}}=30$. Increasing $m$ has the more significant effect on improving convergence than increasing $l_{\rm{max}}$. We also found that increasing $l_{\rm{max}}$ has a greater impact on the rate of convergence than increasing $n_{\rm{max}}$. This informed our decision to set $l_{\text{max}}=40$ and $n_{\text{max}}=30$, and, following [16], we set $m=6$. These choices proved sufficient for yielding an RSET satisfying the conservation equation
@@
From Fig. 3, we see that the scalar condensate is regular everywhere outside the event horizon (see the analysis in Sec. V.3 below for its behaviour on the horizon) and monotonically decreasing as the radial coordinate $r$ increases. The scalar condensates for massless [32] and massive [33] scalar fields on a Schwarzschild black hole are similarly monotonically decreasing as $r$ increases. For a neutral scalar field on a Reissner-Nordström black hole [34], again the scalar condensate is monotonically decreases as $r$ increases, except when the charge of the black hole is close to its maximum value $Q=M$, which is not the situation we are considering here. We also note that in Fig. 3, the scalar condensate is negative except for a region close to the event horizon, whereas the scalar condensate is positive everywhere outside the event horizon for the neutral scalar field results in Refs. [32, 33, 34] (however, this is not the case for a neutral scalar field on a higher-dimensional black hole [23, 24]).
@@ V.3 Expectation values on the event horizon
One such approach is that of [36, 37], where uniform approximations for the radial solution $q_{nl}(r)$ were used to calculate the event horizon values of the scalar condensate and the renormalized stress energy tensor for a field in the Hartle-Hawking state. Even though this method was developed for the case of a neutral scalar field, the results for any mode sums involving the $n=0$ mode $q_{0l}(r)$, such as (V.3, 87), carry through to the charged case with only minor adjustments, leading to the following quasi-analytical expression for the scalar condensate on the event horizon:
@@
In Fig. 6 we compare the results of our numerical computations from the previous section (yellow dots) with the on-horizon expression (V.3), shown as a blue dot. The value obtained by extrapolating the off-horizon results to the horizon agrees with the quasi-analytical expression (V.3) to 9 significant figures.
@@
Inserting Eqs. (V.3) and (B) into the above expression and evaluating for the parameter set considered in this paper, gives the result shown in Fig. 7. We find agreement with the off-horizon numerical results extrapolated to the horizon to 9 significant figures.
Figure 7: Nonvanishing RSET components $\langle{\hat{T}}^{t}{}_{t}\rangle$ (top), $\langle{\hat{T}}^{r}{}_{r}\rangle$ (middle) and $\langle{\hat{T}}^{\theta}{}_{\theta}\rangle$ (bottom) near the black hole event horizon. Yellow dots are computed numerically using the method in Sec. V.2, while blue dots are the on-horizon results (90, 91).
@@
As for the scalar condensate, we may calculate the derivative of the current by exploiting our knowledge of the exact horizon value to obtain the derivative there numerically, without recourse to evaluating any sums involving the $n=1$ mode. The result of this computation is shown in Fig. 8. We find agreement with the extrapolated horizon values to 7 significant figures.
@@ Acknowledgements.
The authors wish to thank Eoin Scanlon for helpful conversations at the early stages of this project. The work of E.W. is supported by STFC grant number ST/X000621/1. E.W. also acknowledges support of the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université), and LabEx CARMIN (ANR-10-LABX-59-01). This work makes use of the Black Hole Perturbation Toolkit [30]. G.M. and E.W. acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing.

### v2

@@ I Introduction
The development, over the past ten years, of two new methods for computing renormalized expectation values has greatly facilitated these calculations and enabled the properties of a neutral quantum scalar field on a much wider range of black hole space-times to be explored. The “pragmatic mode-sum” method [13, 14] works on the original Lorentzian space-time and is notable for its applicability to rotating as well as static black holes [15]. In contrast, the “extended coordinates” method [16] involves the Euclidean space-time. The key feature of the latter approach is that the Hadamard parametrix to any order in the coordinate separation can be expressed as a mode-sum, resulting in a renormalized Green function that converges to any order we desire; the rate of convergence being directly linked to the order of the parametrix we subtract. This method also has the advantage that it is approximately uniform, including close to the horizon.
@@ II Euclidean Green function
$G_{\textrm{\tiny{E}}}(x,x^{\prime})=G^{\textrm{\tiny{(S)}}}(x,x^{\prime})+G^{\textrm{\tiny{(A)}}}(x,x^{\prime}),$ (14a)
@@ III Expansion of the Hadamard Parametrix
and $\Delta r=r-r^{\prime}$ is the separation of the points in the radial direction. The coefficients in the expansions (22) can then be found by substituting (22) into the relevant equations (18, 21) and solving order by order, treating $\varpi\sim s\sim\Delta r\sim\mathcal{O}(\epsilon)$. We begin by substituting the ansatz (22a) into the defining relation
$2\sigma=\sigma_{\alpha}\sigma^{\alpha},$ (24)
writing the derivatives of the extended coordinates (23) as series in $(\varpi,s,\Delta r)$. Solving order by order in $\epsilon$ gives the coefficients $\sigma_{abc}(r)$, which are unchanged from those arising in the neutral case [23, 24, 16]. This expansion for $\sigma(x,x^{\prime})$ is then used, together with (18), to find the coefficients in $U(x,x^{\prime})$ (22b). Next, having the expansions of $\sigma(x,x^{\prime})$ and $U(x,x^{\prime})$, we use these in (21b) to find the coefficients in the expansion of $V_{0}(x,x^{\prime})$, and then, having $V_{k}(x,x^{\prime})$, we can use (21a) to find the coefficients in the expansion of $V_{k+1}(x,x^{\prime})$. Finally, the resulting expansions (22) are combined into the terms appearing in (16). The algebraic manipulations required to find the expansions (22) to the order required are extremely lengthy, and are therefore performed in Mathematica. Further details can be found in [23, 24, 16].
A key feature of this particular choice of “extended coordinates” is that in the limit $\Delta r\to 0$, we have $s^{2}=2\sigma+\mathcal{O}(\epsilon^{3})$. Note also that, although the Hadamard parametrix itself (16) is independent of the quantum state, this particular choice of expansion parameters (23) is adapted specifically to quantum fields in thermal states. This is easily seen in the fact that the extended time coordinate $\varpi$ (23a) is periodic with periodicity related to $\kappa$. This is an essential feature of the method, since the goal is to obtain a mode-sum representation of the parametrix (16) that has the same form as the Euclidean Green function (II) so that a mode-by-mode subtraction can be performed.
@@
Let us deal first with the direct part of the parametrix, $U(x,x^{\prime})/\sigma(x,x^{\prime})$. If we expand the direct part in our extended coordinates up to order $\mathcal{O}(\epsilon^{2m})$ and then take the partial coincidence limit $\Delta r=0$, we obtain
@@
The coefficients are now labelled by a $\mathcal{T}$ to indicate these come from the tail part of the parametrix and the superscript $(\mathrm{L})$ indicates coefficients of terms that contain a logarithm. Again, the first few of these coefficients are listed in App. A.
@@ IV.1 Euclidean vs Lorentzian expressions
An immediate difficulty arises when one comes to computing renormalized expectation values using Euclidean methods: how do we translate the formal Lorentzian expressions for the renormalized expectation values [6] into their Euclidean counterpart? For a neutral scalar field on a static space-time, this is a straightforward matter since one simply replaces the Feynman Green function with the corresponding Euclidean Green function via the mapping
@@
$-iG_{\textrm{\tiny{F}}}(x,x^{\prime})=\langle 0|\mathrm{T}\{\hat{\Phi}(x)\,\hat{\Phi}^{\dagger}(x^{\prime})\}|0\rangle,$ (30)
@@ IV.3 Renormalized current
Now the only nonzero component of the current for a charged scalar field in the Hartle-Hawking state on a static spacetime with an electrostatic potential is the time component, which is
@@
$j_{nl}^{(m)}(r)=$ $\sum_{a=0}^{m-1}\sum_{b=0}^{a}\sum_{k=0}^{m-a}\binom{\tfrac{1}{2}}{k}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{DO}_{ab}^{(\textrm{\tiny{R}})}(r)\Big{[}(2a+2b+1)\Psi_{nl}(a+k,b|r)-(2b+2)f(r)\Psi_{nl}(a+k,b+1|r)\Big{]}$
$+\sum_{a=1}^{m-2}\sum_{b=0}^{a-1}\sum_{k=0}^{m-a-2}\binom{\tfrac{1}{2}}{k}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{TO}_{ab}^{(\textrm{\tiny{R}})}(r)\Big{[}(2a+2b+3)\Psi_{nl}(a+k+1,b|r)$
@@
$+\sum_{a=0}^{m-2}\sum_{b=0}^{a}\sum_{k=0}^{m-a-2}\binom{\tfrac{1}{2}}{k}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{TO}_{ab}^{(\textrm{\tiny{L}})}(r)\Big{[}(2b+1)\chi_{nl}(a,b+k|r)+(2a-2b)f(r)\chi_{nl}(a-1,b+k+1|r)\Big{]}$
$+\sum_{a=0}^{m-2}\sum_{k=0}^{m-a-2}\binom{\tfrac{1}{2}}{k}\left(-\frac{\kappa^{2}}{4}\right)^{k}\mathcal{TO}_{aa}^{(\textrm{\tiny{L}})}(r)\Big{[}2f(r)\Psi_{nl}(a+k+1,0|r)\Big{]}.$ (61b)
@@ IV.4 Renormalized stress energy tensor
using Synge’s rule and the fact that $w$ is independent of time. The last term on the right-hand side we can compute by simply taking a radial derivative of the interpolation function for $w$. The salient point is that now the covariant derivatives in the first term on the right-hand side are at different space-time points and so act as partial derivatives. This can now be computed from taking time derivatives of the mode-sum representation of $W^{\textrm{\tiny{(S)}}}(x,x^{\prime})$ already derived (IV.2). Explicitly, we have
@@
$=$ $~{}\frac{1}{8\pi^{2}}\sum_{l=0}^{\infty}(2l+1)\sum_{n=-\infty}^{\infty}\frac{n^{2}\kappa^{2}}{f(r)}\left\{g_{nl}(r)-k_{nl}^{(m)}(r)\right\}$
@@
$\widetilde{w}^{\phi}{}_{\phi}=\widetilde{w}^{\theta}{}_{\theta}=-\Big{[}\nabla^{\phi^{\prime}}\nabla_{\phi}W^{\textrm{\tiny{(S)}}}(x,x^{\prime})\Big{]}+\frac{f(r)}{2r}w_{,r},$ (73)
@@ V.2 Numerical results
We now use the expressions of the previous section (IV.2, 55, 69) to calculate the renormalized scalar condensate, current and RSET. We consider a Reissner-Nordström black hole with charge $Q=M/2$ and use units in which $M=1$. In these units the event horizon is located at $r=r_{+}=1+{\sqrt{3}}/2\approx 1.86602540$ to 9 s.f. (78). We assume that the scalar field is minimally coupled to the scalar curvature, so that $\xi=0$, and set the scalar field mass to be $\mu M=1/10$, whilst the scalar field charge is $qM=1/4$. With these parameters we have $\mu>\frac{qQ}{r_{+}}$ and therefore no classical charge superradiance [31]. In this case the Euclidean Green function (II) is uniquely defined. The presence of superradiant modes complicates the definition of a Hartle-Hawking-like state for a charged scalar field on a Reissner-Nordström black hole [20] and therefore, the construction of the Euclidean Green function. For this reason, we do not consider the possibility of charge superradiance in this paper, although we expect the framework presented in this paper for implementing Hadamard renormalization will be unaffected by the presence of superradiant modes. Finally, we also set the arbitrary renormalization length scale $L=M=1$ (16).
@@
The next stage is to compute the mode sums appearing in (IV.2, 61a, IV.4, IV.4), performing the sum over $n$ first, before the sum over $l$. As in the neutral case [23, 24], the summand in the sum relevant for calculating the scalar condensate (IV.2), namely $\{\Re\left(g_{nl}(r)\right)-k^{(m)}_{nl}(r)\}$, converges as ${\mathcal{O}}(n^{-2m-3})$ for fixed $l$ and large $n$, as demonstrated in Fig. 1 for $m=0,1,2$.
@@
The number of modes required for final answers for the renormalized expectation values to the desired precision can be reduced by increasing the order $m$ in the expansion of the Hadamard parametrix (IV.2, 61), at the cost of increased computation time of the required regularization parameters (43, 47). Table 1 shows the effect of increasing the expansion order $m$ and $l_{\rm{max}}$ on the convergence of results for the RSET component $\langle\hat{T}^{t}_{~{}t}\rangle$, with $n_{\rm{max}}$ fixed to be $n_{\rm{max}}=30$. Increasing $m$ has the more significant effect on improving convergence than increasing $l_{\rm{max}}$. We also found that increasing $l_{\rm{max}}$ has a greater impact on the rate of convergence than increasing $n_{\rm{max}}$. This informed our decision to set $l_{\text{max}}=40$ and $n_{\text{max}}=30$, and, following [16], we set $m=6$. These choices proved sufficient for yielding an RSET satisfying the conservation equation
@@
From Fig. 3, we see that the scalar condensate is regular everywhere outside the event horizon (see the analysis in Sec. V.3 below for its behaviour on the horizon) and monotonically decreasing as the radial coordinate $r$ increases. The scalar condensates for massless [32] and massive [33] scalar fields on a Schwarzschild black hole are similarly monotonically decreasing as $r$ increases. For a neutral scalar field on a Reissner-Nordström black hole [34], again the scalar condensate monotonically decreases as $r$ increases, except when the charge of the black hole is close to its maximum value $Q=M$, which is not the situation we are considering here. We also note that in Fig. 3, the scalar condensate is negative except for a region close to the event horizon, whereas the scalar condensate is positive everywhere outside the event horizon for the neutral scalar field results in Refs. [32, 33, 34] (however, this is not the case for a neutral scalar field on a higher-dimensional black hole [23, 24]).
@@ V.3 Expectation values on the event horizon
One such approach is that of [36, 37], where uniform approximations for the radial solution $q_{nl}(r)$ were used to calculate the event horizon values of the scalar condensate and the renormalized stress energy tensor for a field in the Hartle-Hawking state. Even though this method was developed for the case of a neutral scalar field, the results for any mode sums involving the $n=0$ mode $q_{0l}(r)$, such as (V.3, 89), carry through to the charged case with only minor adjustments, leading to the following quasi-analytical expression for the scalar condensate on the event horizon:
@@
In Fig. 6 we compare the results of our numerical computations from the previous section (yellow dots) with the on-horizon expression (V.3), shown as a blue dot. The value obtained by extrapolating the off-horizon results to the horizon agrees with the quasi-analytical expression (V.3) to nine significant figures.
@@
Inserting Eqs. (V.3) and (B) into the above expression and evaluating for the parameter set considered in this paper, gives the result shown in Fig. 7. We find agreement with the off-horizon numerical results extrapolated to the horizon to nine significant figures.
Figure 7: Nonvanishing RSET components $\langle{\hat{T}}^{t}{}_{t}\rangle$ (top), $\langle{\hat{T}}^{r}{}_{r}\rangle$ (middle) and $\langle{\hat{T}}^{\theta}{}_{\theta}\rangle$ (bottom) near the black hole event horizon. Yellow dots are computed numerically using the method in Sec. V.2, while blue dots are the on-horizon results (92, 93).
@@
As for the scalar condensate, we may calculate the derivative of the current by exploiting our knowledge of the exact horizon value to obtain the derivative there numerically, without recourse to evaluating any sums involving the $n=1$ mode. The result of this computation is shown in Fig. 8. We find agreement with the extrapolated horizon values to seven significant figures.
@@ Acknowledgements.
The authors wish to thank Eoin Scanlon for helpful conversations at the early stages of this project. The work of E.W. is supported by STFC grant number ST/X000621/1. E.W. also acknowledges support of the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université), and LabEx CARMIN (ANR-10-LABX-59-01). This work makes use of the Black Hole Perturbation Toolkit [30]. G.M. and E.W. acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing. Data supporting this publication can be freely downloaded from the University of Sheffield research data repository at [54], under the terms of the Creative Commons Attribution (CC-BY) licence.